# Lightweight Embedded FPGA Deployment of Learned Image Compression with Knowledge Distillation and Hybrid Quantization

### v3

@@ Abstract
Keywords— FPGA, Deep Learning, Learned Image Compression, Autoencoder, Quantization, Pruning, Knowledge Distillation.
@@ I Introduction
Learned Image Compression (LIC) [1] replaces traditional coding tools in standardized codecs with an end-to-end learnable autoencoder-based model. First, the image is projected into a low-dimensional latent space by the encoder. This representation is then quantized and entropy-coded into a binary bitstream. At the receiver end, the bitstream is entropy-decoded, and the decoder recovers an approximation of the original image from the latent representation. While adding uniform noise during training helps cope with the non-differentiability of quantization, the Generalized Divisive Normalization (GDN) activation minimizes mutual information in the latent space, improving RD efficiency. To date, LIC outperforms early still-image standards such as JPEG and JPEG2000 and performs on par with the intra profile of recent standards such as H.266/VVC.
@@
State-of-the-art hardware LIC implementations like [20, 21, 22] are highly hardware-efficient and can achieve real-time latency (in the case of [22]) by leveraging advanced design techniques and optimizing DSP efficiency. However, such design techniques imply a deep understanding of the specific hardware platform the implementation is designed for and a time-consuming exploration of the design space to meet a latency deadline. Also, real-time operations are often achieved at the expense of RD performance, resulting in significant drops when the hardware implementation is compared to the original LIC model it is based upon. Such drawbacks have prompted our research, where we aim to balance hardware and software optimization by adopting design choices at all levels while ensuring that real-time hardware operations do not come at the expense of RD efficiency by designing ad-hoc components.
The main contributions of this work can be summarized as follows.
@@
  • Latency model: We analytically model the relationship between the number of convolutional channels of a LIC architecture and its latency in terms of FPS. Showing that computational complexity dominates this relationship, we derive an upper bound to latency that is sufficient to dimension the convolutional layers’ width to meet a target frame rate.
@@
  • Knowledge Distillation: We design a framework for distilling a reference LIC teacher model into a smaller student with fewer convolutional channels while keeping the RD efficiency close to the reference. To the best of our knowledge, our work is the first to leverage knowledge distillation to train smaller LIC models suitable for embedded deployment.
@@
  • GDN Implementation: We introduce an implementation of the Generalized Divisive Normalization layer that is hardware-friendly yet preserves RD efficiency when the LIC model is deployed on FPGA. This is a major departure from the approach by Sun et al. [22], where the GDN is replaced with a simpler ReLU to improve latency.
@@
  • Hybrid Quantization-Aware Training (QAT) and Pruning: We introduce a hybrid QAT-based pruning method, which outperforms Post-Training Quantization (PTQ) in terms of RD efficiency. This approach, combined with mixed precision (MP) for the GDN layers, enhances the model’s RD performance while maintaining hardware efficiency. We also propose a structured pruning approach that optimizes model size and latency, contributing to better resource utilization and lower latency on FPGA.
@@
  • Finally, we present a fully pipelined architecture that maximizes hardware resource utilization.
@@
Our experiments show state-of-the-art results in terms of latency, energy efficiency, and resource utilization on the ZCU102 FPGA board (Fig.9 and Table IV). We also outperform state-of-the-art competitors in terms of RD efficiency, i.e., we enable video quality better than other hardware LIC methods despite the better latency. Most importantly, our approach allows meeting different target frame rates without requiring a complex exploration of the hardware design space.
@@
The outline of this paper is as follows. Section II introduces the required background on LIC and discusses the limits of existing hardware implementations. Section III introduces our proposed design methodology involving knowledge distillation, pruning, quantization, and hardware pipelining. Finally, Section IV provides experimental results for the hardware-implemented models on the ZCU102 boards in terms of RD efficiency, latency, power consumption, and resource utilization. In Section IV-D, an in-depth ablation study of each key component of the proposed architecture quantifies the contribution of each component to our overall gains.
Figure 1: Overall proposed workflow for embedded LIC
@@ II Background and Related Works
In this section, we first provide the relevant background on Learned Image Compression (LIC). Then, we overview the existing approaches to hardware friendly LIC, highlighting existing limitations that motivate our work.
@@ II-A A primer on Learned Image Compression
Learned Image Compression (LIC) has demonstrated the ability to match or even outperform standardized codecs in terms of rate-distortion (RD) performance. Foundational works, such as those by Ballé et al [2] and Theis et al. [23] leveraged a simpler convolutional autoencoder architecture. At the transmitter side, a convolutional encoder $g_{a}$ preliminary compresses the image $x$ projecting it onto a low-dimensional hidden latent space $y$ (nonlinear analysis coding). The latent space elements are quantized adding uniform noise at learning time to mimic the rounding operation as a proxy of the non differentiable quantization. The quantized hidden space $\tilde{y}$ is then entropy coded and put on the wire as a binary bitstream. In early works, the probability of each $i$-th element $\tilde{y}_{i}$ of the latent space is learned at training time (fully factorized) without any assumption on the underlying distribution shape (non-parametric). At the receiver side, the bistream is entropy-decoder and a convolutional decoder $g_{s}$ recovers an approximation $\tilde{x}$ of the original image (nonlinear synthesis).
The model is optimized end-to-end by backpropagating and descending the error gradient to minimize the cost function
@@
where $\lambda$ adjusts the trade-off between the rate of the compressed bitstream and the distortion $D$ of the reconstructed image. The rate $R$ of the latent space is estimated by an auxiliary ad-hoc neural network. The distortion between the original image $x$ and the decoded image $\tilde{x}$ is measured either in Mean Squared Error (MSE) terms or perceptually via the Structural Similarity Index (SSIM).
The model relies on the Generalized Divisive Normalization (GDN) [24] activation at the encoder side (and relative inverse GDN at the decoder side). The GDN activation Gaussianizes the latent space $y$ reducing the mutual information amongits components, improving the compression efficiency over the ubiquitous ReLU-like activations used for hidden layers [9]. Such scheme outperformed the JPEG and JPEG2000 standards despite the relatively simple entropy model it relies upon.
The above scheme was improved in [3] introducing a secondary latent space known as the hyperprior. Rather than learning the distributions of the latents $y$ at training time, they are now predicted for each encoded image exploiting the fact that the scale $\theta$ of neighbor components usually varies jointly. The hyperprior latent representation $z$ is first quantized using a factorized entropy model and delivered to the receiver within the bitstream. The main quantized latent space $\tilde{y}$ is then entropy coded assuming a Gaussian distribution with $\mu=0$ and using the scales recovered by the hyperprior decoder available also to the receiver. The model can still be trained end-to-end yet simply adding to the cost function the estimated entropy of the hyperprior. Despite the burden of delivering the compressed hyperprior as side information, such scheme performs en pair with the H.265/HEVC Intra profile (BPG). Clearly, this scheme is more complex due to the presence of hyper encoder and decoder, that are in turn implemented as a nested convolutional autoencoder with GDN normalization.
@@
The works in [25, 7] combined previous architectures with a context-based auto-regressive entropy model to capture more local spatial correlation by exploiting already decoded parts of images. A scheme of this approach is partially presented in Fig. 2, while [26] introduced channels-block decoding and performed block-based encoding to exploit the already decoded part of the latent representation. In more recent works other techniques have been exploited to improve RD performance; [8] replaced simple Gaussian distribution with a mixture of Gaussians and introduced an attention module to enhance entropy estimation. In [27] local attention is exploited to combine the local-aware attention with the global-related feature learning and to date, it remains among the best-performing architectures. We selected this work simply because it represents one of the most recent ones that subsequent works have built upon;
At this point, we note that the existing literature on hardware-friendly implementations has mostly explored CNN-based models due to the difficulty of implementing in HW transformer modules. Additionally, the existing literature has particularly explored [25] as it offers good RD performance and HW efficiency as well and that’s also why in our work we focus on this model rather than on more modern models. At this point we review notable implementations of this model like [20, 22] saying that they achieve realtime operations yet at the expense of RD performance and how, especially the lack of GDN. We conclude that the lack of LIC hardware solution that is both hardware friendly and RD performing motivates our work.
@@
The literature on hardware-friendly image compression (LIC) models is sparse, with existing works predominantly focusing on CNN-based approaches rather than Transformer-based models. This is largely due to the inherent challenges of efficiently implementing Transformer modules in hardware. Among CNN-based models, [22] stands out for building upon [21] and achieving an effective balance between rate-distortion (RD) performance and hardware efficiency, making it the current state of the art in the field. Consequently, our work adopts this model rather than more modern architectures. Notable hardware implementations, such as [22] and [21], have demonstrated real-time processing capabilities but often at the cost of RD performance. These limitations stem from the absence of Generalized Divisive Normalization (GDN) layers and reliance on computationally intensive architectures from the literature. Additionally, [20] targets high-resolution deployment, achieving 4K video decoding, though at frame rates significantly below real-time requirements. The current lack of a hardware solution for LIC that simultaneously delivers hardware efficiency and superior RD performance underscores the motivation for our work, which seeks to develop novel techniques to effectively train and deploy computationally intensive LIC models for real-time applications. In [28]A hierarchical VAE architecture, originally designed for generative image modeling, is utilized for LIC, redefining their probabilistic model to allow easy quantization and practical entropy coding.
TABLE I: Components and Symbols for the Hyperprior LIC model
Component Symbol
Input Image $\mathbf{x}$
Encoder $f(\mathbf{x};\theta_{e})$
Latents $\mathbf{y}$
Latents (quantized) $\hat{\mathbf{y}}$
Latents (quantized, student) $\hat{\mathbf{y}}_{s}$
Latents (quantized, teacher) $\hat{\mathbf{y}}_{t}$
Decoder $g(\hat{\mathbf{y}};\theta_{d})$
Hyper Encoder $f_{h}(\mathbf{y};\theta_{he})$
Hyper-latents $\mathbf{z}$
Hyper-latents (quantized) $\hat{\mathbf{z}}$
Hyper Decoder $g_{h}(\hat{\mathbf{z}};\theta_{hd})$
Context Model $g_{cm}(\mathbf{y}_{<i};\theta_{cm})$
Entropy Parameters $g_{ep}(\cdot;\theta_{ep})$
Reconstruction $\hat{\mathbf{x}}$
Reconstruction (student) $\hat{\mathbf{x}}_{s}$
Reconstruction (teacher) $\hat{\mathbf{x}}_{t}$
@@ III Proposed Method
In this section, we introduce our proposed method consisting in a initial knowledge distillation from a large teacher to a smaller student. Next, the student is implemented using a hardware-friendly GDN/iGDN implementation, the model is pruned, quantized and finally deployed on hardware.
@@ III-A Knowledge Distillation for LIC
This section proposes a Knowledge Distillation (KD) strategy to teach a student LIC model from a larger reference teacher. We first propose a straightforward strategy to dimension a student model with lower complexity than a reference teacher and then we discuss the procedure to distill knowledge from the teacher and into the student.
@@ III-A1 Designing and training a student model
We recall that, given a reference LIC model, our ultimate goal is to deploy it on some hardware platform while meeting some target latency. In approaches like [SUN24], the model complexity is cut by altering its topology (e.g., replacing the GDN activations with RELUs) and via a thorough exploration of the hardware design space. However, altering the model topology badly impacts its RD-efficiency, whereas exploring the hardware design space can be a time-consuming activity. We point out that in models like [3], the main source of complexity lies in the filtering operations performed by the convolutional layers. Multiple hyperparameters drive such complexity, including the number of layers, number of channels per layer, filter sizes, downsampling strategies, etc. Our approach consists in controlling the model complexity by reducing the number of channels per convolutional layer without altering the model topology otherwise. Indeed, exploring one single model hyperparameter is way simpler than exploring the multiple options available in the hardware design space. Also, this approach aligns with established practices in the LIC literature, where reducing channels is commonly used to create smaller, more computationally efficient models [3, 9]. Therefore, we initially design a lower-complexity student model that has a fraction of the number of channels of the reference model. Makes exception to this design rule the dimensions of the last convolutional layer of each encoder and decoder so to preserve the latent spaces dimensions. Preserving the latent spaces dimensions is key for distilling knowledge from the teacher, as discussed below. The student main-model parameters are initialized at random while the student hyperprior model utilizes the teacher’s hyperprior, the student is trained minimizing the same $\mathcal{L}_{\text{initial}}=\mathcal{R}+\lambda\mathcal{D}$ loss function over the same training set. However, such student is less RD-efficient than the reference, as we experimentally show later on. Therefore knowledge must be distilled into the student from the reference model that acts as a teacher, as discussed below.
III-A2 Fune-tuning the student
After designing and preliminarily training the student as above, knowledge is distilled from the teacher to recover the lost RD-efficiency. The dotted boxed in Fig.2 shows how our KD strategy plugs in the training of our reference LIC model. In detail, the student is fine-tuned by minimizing the following loss function
@@ III-A2 Fune-tuning the student
where $\phi_{i}(\mathbf{x})$ denotes the feature map of the reconstructed image $\mathbf{x}$ at layer $i$. The perceptual loss encourages the student model to focus on higher-level semantic similarities between images, rather than merely pixel-wise similarity. Layers 3, 4, and 5 of ResNet-50 are selected as they correspond to increasingly abstract feature representations, progressively moving from edges and textures to object structures. Namely, Layer 3 captures low-level features like edges and textures, essential for preserving fine details in compression. Layer 4 captures mid-level features such as object parts, crucial for maintaining structural coherence. Finally, Layer 5 captures high-level features, representing more abstract and semantic information, essential for preserving the perceptual quality of the image.
Finally, the third term in the loss function is the same RD cost term used to preliminary train the student. During fine-tuning, the three terms of the loss function loss weights $\alpha,\beta,\gamma$ are dynamically adjusted as follows. Early on during fine-tuning, $\alpha$ is set to 1.0 and $\beta$ is set to 0.1 to drive the student latent space to match the teacher latent space and thus achieve similar latent space entropy. Later on, $\alpha$ is reduced to 0.1 and $\beta$ is increased to 1.0 to promote visual quality. The transition between the early and late phases is automated by monitoring the performance plateaus. For instance, once the latent space loss $\mathcal{L}_{\text{latent}}$ stabilizes, the weights are automatically adjusted to emphasize the output space. Empirically, the number of training epochs for each phase can be set to 500 each which meets all the requirements we’ve outlined here on our chosen dataset. The $\gamma_{\text{}}$ weight is set to 0.5 throughout the distillation to maintain balance with the RD loss and avoid catastrophic forgetting of the original end-to-end compression task. Table II in the experimental section provides all the relevant hyperparameters for the distillation process. It is worth mentioning that throughout the training the initially trained teacher hyperprior model is being reused.
TABLE II: Hyperparameters for the KD Training Scheme
Hyperparameter Initial Fine Tuning
Training Early Late
Learning Rate 1e-3 1e-4 1e-6
Decay [Step, Rate] 500k,0.95 250k, 0.9 250k, 0.9
Batch Size [256x256] 16 8 8
$\alpha$ NA 0.1 1.0
$\beta$ NA 1.0 0.1
$\gamma$ NA 0.5 0.5
Steps 5M 2.5M 2.5M
$\lambda$ 0.0016, 0.0032, 0.0075, 0.045

### v5

# Lightweight FPGA Deployment of Learned Image Compression with Knowledge Distillation and Hybrid Quantization

@@ Abstract
Keywords— FPGA, Deep Learning, Learned Image Compression, Autoencoder, Quantization, Pruning, Knowledge Distillation. [^1] This work is carried out under the framework of the NewEmma (Neural nEtwork Watermarking for Energy efficient Mobile Multimedia Applications) project funded by DIGICOSME.
@@ I Introduction
Learned Image Compression (LIC) [1, 2] replaces traditional coding tools in standardized codecs with an end-to-end learnable autoencoder-based model. First, the image is projected into a low-dimensional latent space by the encoder. This representation is then quantized and entropy-coded [3, 4] into a binary bitstream. At the receiver end, the bitstream is entropy-decoded, and the decoder recovers an approximation of the original image from the latent representation. While adding uniform noise during training helps cope with the non-differentiability of quantization, the Generalized Divisive Normalization (GDN) activation minimizes mutual information in the latent space, improving RD efficiency. To date, LIC outperforms early still-image standards such as JPEG and JPEG2000 [5] and performs on par with the intra profile of recent standards such as H.266/VVC [6].
@@
State of the art hardware LIC implementations like [25], [26, 27, 28, 29] are highly hardware-efficient and can achieve real-time latency (in the case of [29, 28]) by leveraging advanced design techniques and optimizing DSP efficiency, although [28] uses the old [7] model resulting in a considerably lower RD performance compared to [29] and [26]. The Currently [29] provides the best system level and RD performance with real-time encoding at reasonable RD quality, however such design techniques imply a deep understanding of the specific hardware platform the implementation is designed for and a time-consuming exploration of the design space to meet a latency deadline. Also, real-time operations are often achieved at the expense of RD performance, resulting in significant drops when the hardware implementation is compared to the original LIC model it is based upon. Such drawbacks have prompted our research, where we aim to balance hardware and software optimization adopting design choices at all levels on one hand and ensuring both real-time hardware operations do not come at the expense of RD-efficiency by designing some ad-hoc components on the other.
The main contributions of this work are:
@@
  • Latency Model: We analytically model the relationship between convolutional channels in a LIC architecture and its latency (FPS). By showing that computational complexity dominates this relationship, we derive an upper bound for latency, enabling the dimensioning of convolutional layer width to meet target frame rates.
@@
  • Knowledge Distillation: We introduce a framework for distilling a reference LIC teacher model into a smaller student model, while maintaining RD-efficiency. To our knowledge, this is the first work leveraging knowledge distillation for training smaller LIC models suitable for embedded deployment.
@@
  • GDN Implementation: We present a hardware-friendly implementation of the Generalized Divisive Normalization (GDN) layer that preserves RD-efficiency on FPGA. Unlike Sun et al. [29, 28, 26], who replace GDN with ReLU for latency improvement, our approach retains GDN’s contribution to RD performance.
@@
  • Hybrid QAT and Pruning: We propose a hybrid QAT method that outperforms Post-Training Quantization (PTQ) in RD efficiency. Coupled with mixed precision for GDN layers, this enhances RD performance while maintaining hardware efficiency. Additionally, structured pruning optimizes model size and latency, improving FPGA resource utilization.
@@
  • Pipelined Architecture: We present a fully pipelined architecture that maximizes hardware resource utilization for through parallel patching of high resolutions inputs.
@@
Our experiments show state-of-the-art results in terms of latency, energy efficiency and resource utilization on the ZCU102 FPGA board (Fig.9 and Table VII). We also outperform state-of-the-art competitors in RD-efficiency, achieving higher video quality than other hardware LIC methods while maintaining lower latency. Most important, our approach allows meeting different target frame rates without a complex exploration of the hardware design space.
@@
The outline of this paper is as follows: Section LABEL:sec:background_related provides background on LIC and reviews the limitations of existing hardware implementations. Section III presents our design methodology, including knowledge distillation, pruning, quantization, and hardware pipelining. Finally, Section 4 presents experimental results for hardware-implemented models on the ZCU102 boards, covering RD efficiency, latency, power consumption, and resource utilization. An ablation study analyzes the contribution of each key component to the overall performance gains.
Figure 1: The proposed workflow for training, optimizing, distilling, and deploying LIC models on hardware is comprehensive yet abstracts hardware-specific compilation through Xilinx VITIS-AI APIs
@@ II Background and Related Works
This section first covers the background on Learned Image Compression (LIC) and then reviews existing hardware-friendly LIC approaches, highlighting the limitations that motivate our work.
@@ II-A A primer on Learned Image Compression
Learned Image Compression (LIC) has shown the ability to match or surpass standardized codecs in rate-distortion (RD) performance. Early works by Ballé et al. [7] and Theis et al. [30] used simpler convolutional autoencoders. At the transmitter, a convolutional encoder $g_{a}$ compresses the image $x$ into a low-dimensional latent space $y$ (nonlinear analysis coding). The latent elements are quantized with uniform noise added during training to simulate rounding, acting as a proxy for non-differentiable quantization. The quantized space $\tilde{y}$ is entropy-coded and transmitted as a binary bitstream. In early works, each element $\tilde{y}_{i}$ of the latent space is learned with a non-parametric, fully factorized approach, without assumptions about the underlying distribution shape. At the receiver, the bitstream is decoded, and a convolutional decoder $g_{s}$ reconstructs the approximation $\tilde{x}$ of the original image (nonlinear synthesis).
The model is optimized end-to-end by backpropagating the error gradient to minimize the cost function:
@@
where $\lambda$ controls the trade-off between the rate $R$ of the compressed bitstream and the distortion $D$ of the reconstructed image. The rate $R$ is estimated by an auxiliary neural network, while distortion $D$ is measured using either Mean Squared Error (MSE) or Structural Similarity Index (SSIM). The encoder uses Generalized Divisive Normalization (GDN) [31] activation, with inverse GDN at the decoder. GDN Gaussianizes the latent space $y$, reducing mutual information and improving compression efficiency compared to ReLU-like activations [14]. This approach outperforms JPEG and JPEG2000 standards, despite using a relatively simple entropy model.
The above scheme was improved in [8] introducing a secondary latent space known as the hyperprior. Rather than learning the distributions of the latents $y$ at training time, they are now predicted for each encoded image exploiting the fact that the scale $\theta$ of neighbor components usually varies jointly. The hyperprior latent representation $z$ is first quantized using a factorized entropy model and delivered to the receiver within the bitstream. The main quantized latent space $\tilde{y}$ is then entropy coded assuming a Gaussian distribution with $\mu=0$ and using the scales recovered by the hyperprior decoder available also to the receiver. The model can still be trained end-to-end yet simply adding to the cost function the estimated entropy of the hyperprior. Despite the burden of delivering the compressed hyperprior as side information, such scheme performs on par with the H.265/HEVC Intra profile (BPG). Clearly, this scheme is more complex due to the presence of hyper encoder and decoder, that are in turn implemented as a nested convolutional autoencoder with GDN normalization.
@@
[32, 12] combined previous architectures with a context-based auto-regressive entropy model to capture local spatial correlations by exploiting already decoded image parts, An overview of this approach is shown in Fig. 2. While this improves rate performance, it introduces significant latency due to sequential decoding. The model requires conditioning each pixel on previously decoded ones, limiting parallelism, which is crucial for high throughput on hardware accelerators like FPGAs. Achieving real-time compression would require architectural modifications, such as pipelining or parallel decoding techniques, to reduce latency, as shown in [26].
In [33] local attention is exploited to combine the local-aware attention with the global-related feature learning and to the present date is among the best-performing architectures.
@@
We note that recent hardware-friendly implementations of Learned Image Compression (LIC) have primarily focused on CNN-based models due to the challenges of implementing transformer modules in hardware. Among these, [32] has been widely explored for its strong rate-distortion (RD) performance and hardware efficiency, which is why our work focuses on this model rather than more modern alternatives. We review notable implementations, such as [28] and [29], which achieve real-time operations, but at the cost of RD performance, particularly due to the lack of Generalized Divisive Normalization (GDN) or the use of an older architecture. This gap in hardware solutions that balance both RD performance and hardware friendliness motivates our work. The literature on hardware-friendly image compression (LIC) models is sparse, with existing works predominantly focusing on CNN-based approaches rather than Transformer-based models. This is largely due to the inherent challenges of efficiently implementing Transformer modules in hardware. Among CNN-based models, [29] stands out for building upon [27] and achieving an effective balance between rate-distortion (RD) performance and hardware efficiency, making it the current state of the art in the field. Consequently, our work adopts this model rather than more modern architectures. Notable hardware implementations, such as [29] and [27], have demonstrated real-time processing capabilities but often at the cost of RD performance. These limitations stem from the absence of Generalized Divisive Normalization (GDN) layers and reliance on computationally intensive architectures from the literature. Additionally, [26] targets high-resolution deployment, achieving 4K video decoding, though at frame rates significantly below real-time requirements. The current lack of a hardware solution for LIC that simultaneously delivers hardware efficiency and superior RD performance underscores the motivation for our work, which seeks to develop novel techniques to effectively train and deploy computationally intensive LIC models for real-time applications.
@@ III Proposed Method
In this section, we present our proposed method, which begins with knowledge distillation from a large teacher model to a smaller student. The student model is then implemented using a hardware-friendly GDN/iGDN, followed by pruning, mixed-precision quantization, and final deployment on hardware.
@@ III-A Knowledge Distillation for LIC
This section proposes a novel Knowledge Distillation (KD) scheme to transfer knowledge from a larger reference teacher model to a smaller student LIC model. We first outline a strategy to design a student model with lower complexity than the teacher, followed by the procedure to distill knowledge from the teacher into the student.
@@ III-A1 Designing and training a student model
We recall that, given a reference LIC model, our ultimate goal is to deploy a reference LIC model on a hardware platform while meeting a target latency. In approaches like [29], model complexity is reduced by altering the topology (e.g., replacing GDN activations with ReLUs) and extensively exploring the hardware design space. However, modifying the model topology often leads to a significant loss in RD efficiency, while exploring the hardware design space can be time-consuming, resource-intensive, and highly demanding in terms of hardware expertise. In models like [8], the main source of complexity stems from the filtering operations performed by the convolutional layers. Multiple hyperparameters contribute to this complexity, including the number of layers, channels per layer, filter sizes, down-sampling strategies, etc. Our approach controls model complexity by reducing the number of channels per convolutional layer while keeping the overall model topology unchanged. Exploring a single hyperparameter is far simpler than navigating the extensive hardware design space. Moreover, this approach aligns with established practices in the LIC literature, where reducing channels is a common strategy to create smaller, more computationally efficient models [8, 14]. Therefore, we begin by designing a lower-complexity student model with a fraction of the channels of the reference model. The only exception to this rule is the last convolutional layer of each encoder and decoder, where we preserve the latent space dimensions to ensure effective knowledge distillation from the teacher, as discussed below. the student model’s main parameters are initialized randomly, while its hyperprior model uses the teacher’s hyperprior. The student is trained by minimizing the same loss function, $\mathcal{L}_{\text{initial}}=\mathcal{R}+\lambda\mathcal{D}$, over the same training set. However, as we show experimentally later, this student model is less RD-efficient than the reference. Therefore, knowledge distillation from the teacher model into the student is necessary to guide the student to learn a more optimized RD representation.
III-A2 Fine-tuning the student through output and latent space knowledge distillation
After designing and preliminarily training the student model, knowledge is distilled from the teacher to recover the lost RD-efficiency. The dotted purple boxes in Fig.2 illustrates how our KD strategy integrates into the training of the reference LIC model. Specifically, the student is fine-tuned by minimizing the following loss function:
@@ III-A2 Fine-tuning the student through output and latent space knowledge distillation
where $\phi_{i}(\mathbf{x})$ denotes the feature map of the reconstructed image $\mathbf{x}$ at layer $i$. The perceptual loss encourages the student model to focus on higher-level semantic similarities between images, rather than merely pixel-wise similarity. Layers 3, 4, and 5 of ResNet-50 are selected as they correspond to increasingly abstract feature representations, progressively moving from edges and textures to object structures. Namely, Layer 3 captures low-level features like edges and textures, essential for preserving fine details in compression. Layer 4 captures mid-level features such as object parts, crucial for maintaining structural coherence. Finally, Layer 5 captures high-level features, representing more abstract and semantic information, essential for preserving the perceptual quality of the image. Finally, the third term in the loss function is the same RD cost term used to preliminary train the student. During fine-tuning, the three terms of the loss function loss weights $\alpha,\beta,\gamma$ are dynamically adjusted as described in IV-A1.